OpenStudy (anonymous):

Find the distance between the points (-5, 4) and (3, 1) in simplest radical form

OpenStudy (anonymous):

use the distance formula

OpenStudy (anonymous):

but i dont know how to do radicals

OpenStudy (lgbasallote):

distance formmula is \[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] if you substitute your values you get \[\sqrt{(3 + 5)^2 + (1 - 4)^2}\]

OpenStudy (anonymous):

the answer i got with distance formula was 8.54400374531753

OpenStudy (lgbasallote):

yup that's right

OpenStudy (anonymous):

but it says radical form i dont know how to do that

OpenStudy (lgbasallote):

just simplify \[\sqrt{(3+5)^2 + (1-4)^2}\] then

OpenStudy (lgbasallote):

3+ 5 is 8...8 squared is 64

OpenStudy (lgbasallote):

1 - 4 is -3...-3 squared is 9 so we have \[\sqrt{64 + 9}\]

OpenStudy (lgbasallote):

therefore radical form is \[\sqrt{73}\] got it?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

thanks